OpenStudy (anonymous):

How to integrate to int ( 1/(e ^(3y) -1 ) dy

OpenStudy (zzr0ck3r):

that would be e^(3y) what is the anti derivative of e^(3y)?

OpenStudy (zzr0ck3r):

does that say 1/((e^(3y)) - 1)?

OpenStudy (experimentx):

\[ 1 + \frac{1}{e^{3y} - 1} = \frac{e^{3y}}{e^{3y} - 1}\]

OpenStudy (anonymous):

\[\int\limits ( 1 / (e^(3y) -1 ) dy \]

OpenStudy (anonymous):

basically int ( e^ (3y) - 1) -1 dy

OpenStudy (anonymous):

any solution , I really have forgotten calculus .

OpenStudy (experimentx):

can you integrate the right hand side of the above expression?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The full equation is (dy/dx) = x^2(e^3y - 1)

OpenStudy (experimentx):

huh?? that's a Differential equation

OpenStudy (zzr0ck3r):

these are not the same questions:)

zepdrix (zepdrix):

So you started with this, \[\large \frac{dy}{dx}=x^2(e^{3y}-1)\] You separated variables and got stuck here on this left side?\[\large \int\limits \frac{dy}{e^{3y}-1}=\int\limits x^2 dx\] Oh.. you're offline -_-